# Understanding Domain-Size Generalization in Markov Logic Networks

### v1

@@ 1 Introduction
Given the magnitude and ever-increasing nature of relational data, like social networks and epidemiology data, only a subsample of the data is ever observed. Statistical Relational Learning (SRL) [6, 5] methods integrate logic and probability to learn and infer over such data. However, are parameters estimated from subsampled data a good fit for the model of the larger relational structure? Cosma and Shalizi [22] showed that, for most non-trivial probabilistic models on relations structures, it is probabilistically inconsistent to apply the same model both to the whole relational structure and to its substructures. Jaeger and Schulte [9] extend this analysis to a vast array of SRL models. These results show that, unlike independent and identically distributed (iid) data, relational data does not admit consistency of parameter estimation. That is, it is not true that the maximum likelihood (ML) parameter estimate converges to the true model parameters as the size of the observed data grows. In fact, these results show that the notion of a single true parameter, for relational structures of all sizes, is ill-defined for SRL models.
Lack of probabilistic consistency means that using an SRL model learned on a fixed domain for inference on a domain of different size may lead to poor results. The poor generalization behavior of SRL models across domain sizes is indeed observed in multiple empirical studies [10, 16, 19, 29]. Such issues can be ameliorated by using projective models — probabilistic models where the same parameters can be used for both the whole relational structure and its subsample. Formally, projective models capture probability distributions on relational structures (resp. graphs) of size $n$, where the marginal distribution over substructures (resp. subgraphs) of size $m<n$ does not depend on $n$. However, Cosma and Shalizi [22] also show that no projective model can express probability distributions with complex sufficient statistics, like $k$-cliques for any $k$ larger than two. These results also exclude the possibility of constructing any SRL model with practically desirable First-Order Logic (FOL) features such as transitivity. Given these results, it is unclear what quantitative statements can be made about the generalization behavior of SRL models across domain sizes.
@@ 2 Related work
Lack of probabilistic consistency in probabilistic models on relational structures was first investigated by Cosma and Shalizi [22]. Jaeger and Schulte [8] showed that such issues persist in most practically used SRL models. A large array of works have tried to devise new projective methods [9, 32] or identify and characterize projective fragments of existing SRL models [14, 30], which circumvent these issues. However, most of the proposed new models are currently only of theoretical interest, as no clear way of learning or reasoning with them has been developed. On the other hand, most of the projective fragments of SRL models are rather restrictive.
These theoretical shortcomings are also reflected in the poor generalization behavior of SRL models in practice [10, 16, 19, 29]. Many works provide heuristic solutions [10, 16, 19] for better generalization across domain sizes. A particularly relevant family of formalisms adapts the parameter values with the size of the domain [16, 28]. In formalisms based on directed graphical models, it has been shown that parameter scaling leads to asymptotically projective models [31], but none of the heuristics developed for MLNs are formally motivated in this way [28]. For MLNs, [13] provides a sound approach for estimating the parameters for a larger unseen domain from a smaller subsample of fixed size. However, the practical applicability of this result is unclear.
In comparison to aforementioned works, we analyze the generalization behavior of an MLN in the most natural setting, i.e., the MLN parameters are learned from a subsample of smaller size, and we analyze the behavior of such a distribution on the larger domain. Our analysis theoretically justifies many of the existing heuristic methods [16]. Our results are also relevant to works in ERGMs [21], investigating the relationship between the sample (resp. the substructure for us) and the population (resp. the larger relational structure for us).
@@ 3.1 Basic Definitions
The set of integers $\{1,...,n\}$ is denoted by $[n]$. We use $[m:n]$ to denote the set of integers $\{m,...,n\}$. Wherever the larger set of integers $[n]$ is clear from context, we will use $[\bar{m}]$ to denote the set $[m+1:n]$. For any $d\geq 1$, $\langle n\rangle^{d}$ represents $d$-tuples in $[n]^{d}$, with $d$ distinct elements, appearing in natural order. Hence, $\langle n\rangle^{d}$ forms a standardized representation of the set of all $d$-element subsets in $[n]^{d}$.
@@ 4 Learning in Markov Logic
where $N(\phi_{i},\omega)$ is the number of true groundings of $\phi_{i}$ in $\omega$. The normalization constant $Z(n)$ is called the partition function that ensures that $P^{(n)}_{\Phi}$ is a probability distribution. In the following we provide an example of an MLN which models the spread of COVID-19 due to contact among different individuals and the impact of vaccines.
@@
However, given that most MLNs are not projective, the ML estimate in equation (3) is not the same as the ML estimate in (2). As $m$ may be very large, it can be computationally prohibitive to make the ML estimate for the distribution $P^{(n+m)}_{\Phi}$. Furthermore, in many cases it may be hard to know or guess the value of $m$. Hence, our goal would be to analyze the relation between the distributions $P^{(n)}_{\Phi}(\omega)$ and $P^{(n+m)}_{\Phi}\downarrow[n](\omega)$, and use that analysis to subsequently characterize conditions that lead to better ML parameter estimates for $P^{(n+m)}_{\Phi}(\omega)$, or in other words generalize better to larger domains.
5 Markov Logic and Projectivity
In this section, we analyze MLNs and briefly discuss an existing projective fragment of MLNs known as the $\sigma$-determinate MLNs. We present the necessary machinery for our main results in section 6 and create an intuition for what leads to projectivity, and how any MLN can be made closer to being projective.
@@ 5 Markov Logic and Projectivity
In the following two Lemmas, we decompose the contribution of each $k$-substructure towards the weight $w(\omega)$.
@@
$\sum_{(\phi_{i},a_{i})\in\Phi}a_{i}N(\phi_{i},\omega)=\sum_{k\in[d]}\sum_{\mathbf{c}\in\langle n\rangle^{k}}\sum_{(\phi_{i},a_{i})\in\Phi_{k}}a_{i}N(\phi_{i},\omega\downarrow\mathbf{c})\$
@@
$w(\omega)=w(\omega\downarrow[n])\times w(\omega\downarrow[\bar{n}])\times\prod_{k\in[d]}\prod_{\mathbf{c}\in\langle n+m\rangle^{k}/\langle n\rangle^{k}\cup\langle\bar{n}\rangle^{k}}\!\!\!\!w_{k}(\omega\downarrow\mathbf{c})$ (7)
@@
$=\prod_{k\in[d]}\prod_{\mathbf{c}\in\langle n\rangle^{k}}w_{k}(\omega\downarrow\mathbf{c})\prod_{k\in[d]}\prod_{\mathbf{c}\in\langle m\rangle^{k}}w_{k}(\omega\downarrow\mathbf{c})\prod_{k\in[d]}\prod_{\mathbf{c}\in\langle n+m\rangle^{k}/\langle n\rangle^{k}\cup\langle\bar{n}\rangle^{k}}\!\!\!\!\!\!\!\!w_{k}(\omega\downarrow\mathbf{c})$
$=w(\omega\downarrow[n])\times w(\omega\downarrow[\bar{n}])\times\prod_{k\in[d]}\prod_{\mathbf{c}\in\langle n+m\rangle^{k}/\langle n\rangle^{k}\cup\langle\bar{n}\rangle^{k}}\!\!\!\!w_{k}(\omega\downarrow\mathbf{c})$
@@
$\prod_{\mathbf{c}\in\langle n+m\rangle^{k}/\langle n\rangle^{k}\cup\langle\bar{n}\rangle^{k}}\!\!\!\!w_{k}(\omega\downarrow\mathbf{c})$ (8)
@@
We now discuss an existing projective fragment of MLNs, namely $\sigma$-determinate MLNs. A Markov Logic Network $\Phi:=\{\phi_{i},a_{i}\}_{i}$ is $\sigma$-determinate if its formulas $\phi_{i}$ satisfy that any two atoms appearing in $\phi_{i}$ contain the same variables.
Proposition 1 ( $\sigma$- determinate MLN[9, 23])
A Markov Logic Network $\Phi:=\{\phi_{i},a_{i}\}_{i}$ is projective if its formulas $\phi_{i}$ satisfy that any two atoms appearing in $\phi_{i}$ contain the same variables.
Example 3
Following is an example of a $\sigma$-determinate MLN:
$a_{1}$ $\quad\mathtt{Covid}(x)$
$a_{2}$ $\quad\mathtt{Contact}(x,y)\land\mathtt{Contact}(y,x)$
Example 3 shows a $\sigma$-determinate MLN. Even simple MLNs, such as the one presented in Example 2, can not be represented as a $\sigma$-determinate MLN.
@@ 6 Domain-Size Generalization
The statement follows from equation (7) in Lemma 2. The upper bound is obtained by replacing the multiplicative weight contribution of each tuple in $\langle n+m\rangle^{k}/\langle n\rangle^{k}\cup\langle\bar{n}\rangle^{k}$ with $w_{k}^{max}$, for all $k\in[d]$. And the lower bound is obtained by replacing the weight contribution of all such tuples with $w_{k}^{min}$. ∎
@@
∎As the proof of the lower bound follows analogous to the proof of the upper bound, we defer it to the appendix.
@@
Given a Markov Logic Network $\Phi$, then the following inequality holds:
$\frac{M_{min}}{M_{max}}P_{\Phi}^{(n)}(\omega^{\prime})\leq P_{\Phi}^{(n+m)}\downarrow[n](\omega^{\prime})\leq\frac{M_{max}}{M_{min}}P_{\Phi}^{(n)}(\omega^{\prime})$ (14)
@@
$=\frac{1}{Z(m)M_{min}C_{n,m}}w(\omega^{\prime})\sum_{\begin{subarray}{c}\omega\in\Omega^{(n+m)}\\ \omega\downarrow[n]=\omega^{\prime}\end{subarray}}\frac{w(\omega\downarrow[\bar{n}])M_{max}}{Z(n)}$
$=\frac{1}{Z(m)M_{min}C_{n,m}}w(\omega^{\prime})\frac{Z(m)C_{n,m}M_{max}}{Z(n)}$
@@
$=\log\Delta$
@@ 7 Experiments
In this section, we evaluate the effect of reducing parameter variance on generalization behavior Proposition 2 bounds $w(\omega)$ w.r.t. the maxima and minima of $w_{k}$. However, for almost all worlds, this bound is loose. This is because, for most worlds, not all $k$-tuples chosen from across the domains will have the extreme weights. Note that in general, our goal is to minimize the impact of the term presented in equation (8). Therefore, it is more effective to reduce the spread between all the weights, rather than merely scaling the upper and the lower bound. Also note that for most MLNs, for some $\omega\in\Omega^{k}$, we will have that $w_{k}(\omega)=1$, i.e., none of the formulas in the MLN will be realized on $\omega$. Thus, in most practical cases, to reduce the spread of the weights $a_{i}$, one should reduce their spread around 00.
Multiple approaches discussed in the literature, directly or indirectly, minimize the parameter variance [7, 16]. We empirically evaluate the effects of three such approaches: L1 regularization, L2 regularization, and Domain-Size Aware Markov Logic Networks (DA-MLNs) [16]. Both L1 and L2 regularization directly work to reduce the spread of the parameters: L1 regularization penalizes the sum of the absolute weight values and L2 regularization penalizes the sum of squared weights. In our setting, we only penalize formulas of arity $>1$, because unary formulas do not affect the connecting term discussed in equation (8). A DA-MLN is an adaptation of a regular MLN that reduces the variance of the parameters by down-scaling formula weights depending on the domain size of the dataset it should generalize to. In this section, we will call such datasets target sets. A DA-MLN is then given as follows:
@@ 7.1 Datasets
To provide a thorough analysis of the effects of different methods for generalizing across different domain sizes, we use four datasets commonly used in related literature: Friends & Smokers (FS) [24], IMDB [^3] Dataset available on the Alchemy website: https://alchemy.cs.washington.edu/data/ [15], WebKB\@footnotemark [15] and Nations\@footnotemark [20].
Friends & Smokers (FS). This synthetic dataset captures information about smoking habits, friendships, and cancer diagnoses of a set of people. The data is created by first randomly selecting 40% of a population to be smokers. Then, 30% of the smokers and 10% of the non-smokers are chosen to suffer from cancer. Lastly, friendships are assigned based on smoking habits, with a $0.8$ probability for friendships between people with the same smoking habit, and a $0.1$ probability of friendships between people with different smoking habits. For our experiments, we generate a target dataset of size 500.

### v3

@@ 1 Introduction
Given the magnitude and ever-increasing nature of relational data, like social networks and epidemiology data, only a subsample of the data is ever observed. Statistical Relational Learning (SRL) [6, 5] methods integrate logic and probability to learn and infer over such data. However, are parameters estimated from subsampled data a good fit for the model of the larger relational structure? Shalizi et al. [23] showed that, for most non-trivial probabilistic models on relations structures, it is probabilistically inconsistent to apply the same model both to the whole relational structure and to its substructures. Jaeger et al. [10] extend this analysis to a vast array of SRL models. These results show that, unlike independent and identically distributed (iid) data, relational data does not admit consistency of parameter estimation. That is, it is not true that the maximum likelihood (ML) parameter estimate converges to the true model parameters as the size of the observed data grows. In fact, these results show that the notion of a single true parameter, for relational structures of all sizes, is ill-defined for SRL models.
Lack of probabilistic consistency means that using an SRL model learned on a fixed domain for inference on a domain of different size may lead to poor results. The poor generalization behavior of SRL models across domain sizes is indeed observed in multiple empirical studies [11, 17, 20, 30]. Such issues can be ameliorated by using projective models — probabilistic models where the same parameters can be used for both the whole relational structure and its subsample. Formally, projective models capture probability distributions on relational structures (resp. graphs) of size $n$, where the marginal distribution over substructures (resp. subgraphs) of size $m<n$ does not depend on $n$. However, Shalizi et al. [23] also show that no projective model can express probability distributions with complex sufficient statistics, like $k$-cliques for any $k$ larger than two. These results also exclude the possibility of constructing any SRL model with practically desirable First-Order Logic (FOL) features such as transitivity. Given these results, it is unclear what quantitative statements can be made about the generalization behavior of SRL models across domain sizes.
@@ 2 Related work
Lack of probabilistic consistency in probabilistic models on relational structures was first investigated by Shalizi et al. [23]. Jaeger et al. [9] showed that such issues persist in most practically used SRL models. A large array of works have tried to devise new projective methods [10, 29] or identify and characterize projective fragments of existing SRL models [15, 31], which circumvent these issues. However, most of the proposed new models are currently only of theoretical interest, as no clear way of learning or reasoning with them has been developed. On the other hand, most of the projective fragments of SRL models are rather restrictive.
These theoretical shortcomings are also reflected in the poor generalization behavior of SRL models in practice [11, 17, 20, 30]. Many works provide heuristic solutions [11, 17, 20] for better generalization across domain sizes. A particularly relevant family of formalisms adapts the parameter values with the size of the domain [17, 28]. In formalisms based on directed graphical models, it has been shown that parameter scaling leads to asymptotically projective models [32], but none of the heuristics developed for MLNs are formally motivated in this way [28]. For MLNs, [14] provides a sound approach for estimating the parameters for a larger unseen domain from a smaller subsample of fixed size. However, the practical applicability of this result is unclear. Furthermore, results provided in [14] rely on learning on a larger domain. Hence, the computational complexity of learning can be prohibitively large in real-world settings.
In comparison to the aforementioned works, we analyze the generalization behavior of an MLN in the most natural setting, i.e., the MLN parameters are learned from a subsample of smaller size, and we analyze the behavior of such a distribution on the larger domain. Our analysis theoretically justifies many of the existing heuristic methods [17]. Our results are also relevant to works in ERGMs [22], investigating the relationship between the sample (resp. the substructure for us) and the population (resp. the larger relational structure for us).
@@ 3.1 Basic Definitions
The set of integers $\{1,...,n\}$ is denoted by $[n]$. We use $[m:n]$ to denote the set of integers $\{m,...,n\}$. Wherever the larger set of integers $[n]$ is clear from context, we will use $[\bar{m}]$ to denote the set $[m+1:n]$. For any $d\geq 1$, $\langle n\rangle^{d}$ represents $d$-tuples in $[n]^{d}$, with $d$ distinct elements, appearing in natural order. Hence, $\langle n\rangle^{d}$ forms a standardized representation of the set of all $d$-element subsets in $[n]$.
@@ 4 Learning in Markov Logic
where $N(\phi_{i},\omega)$ is the number of true groundings of $\phi_{i}$ in $\omega$. The normalization constant $Z(n)$ is called the partition function that ensures that $P^{(n)}_{\Phi}$ is a probability distribution. In the following, we provide an example of an MLN which models the spread of COVID-19 due to contact among different individuals and the impact of vaccines.
@@
However, given that most MLNs are not projective [10, 23], the ML estimate in equation (3) is not the same as the ML estimate in (2). As $m$ may be very large, it can be computationally prohibitive to make the ML estimate for the distribution $P^{(n+m)}_{\Phi}$. Furthermore, in many cases it may be hard to know or guess the value of $m$. Hence, our goal would be to analyze the relation between the distributions $P^{(n)}_{\Phi}(\omega)$ and $P^{(n+m)}_{\Phi}\downarrow[n](\omega)$, and use that analysis to subsequently characterize conditions that lead to better ML parameter estimates for $P^{(n+m)}_{\Phi}$, or in other words generalize better to larger domains.
Remark 1
Although projective MLNs can easily be obtained, their expressivity is significantly limited. One projective fragment of MLNs is the $\sigma$-determinate MLNs [10, 24]. A Markov Logic Network $\Phi:=\{\phi_{i},a_{i}\}_{i}$ is $\sigma$-determinate if its formulas $\phi_{i}$ satisfy that any two atoms appearing in $\phi_{i}$ contain the same variables.
Example 3
Following is an example of a $\sigma$-determinate MLN:
$a_{1}$ $\quad\mathtt{Covid}(x)$
$a_{2}$ $\quad\mathtt{Contact}(x,y)\land\mathtt{Contact}(y,x)$
Example 3 shows a $\sigma$-determinate MLN. Even simple MLNs, such as the one presented in Example 2, can not be represented as a $\sigma$-determinate MLN. Hence, our goal in this paper can also be framed as to obtain arbitrarily expressive MLNs that are close to being projective.
5 Markov Logic Across Domain Sizes
In this section, we analyze how the weights induced by an MLN distribute over different parts of the domain. We present the necessary machinery for our main results in Section 6 and create an intuition for what leads to projectivity, and how any MLN can be made closer to being projective.
@@ 5 Markov Logic Across Domain Sizes
In the following two Lemmas, we further decompose the contribution of each $k$-substructure towards the weight $w(\omega)$.
@@
$\sum_{(\phi_{i},a_{i})\in\Phi}a_{i}N(\phi_{i},\omega)=\sum_{k\in[d]}\sum_{\mathbf{c}\in\langle n\rangle^{k}}\sum_{(\phi_{i},a_{i})\in\Phi_{k}}a_{i}N(\phi_{i},\omega\downarrow\mathbf{c})$
@@
$w(\omega)=w(\omega\downarrow[n])\times w(\omega\downarrow[\bar{n}])\times\prod_{k\in[d]}\prod_{\mathbf{c}\in\langle n+m\rangle^{k}\setminus\langle n\rangle^{k}\cup\langle\bar{n}\rangle^{k}}\!\!\!\!w_{k}(\omega\downarrow\mathbf{c})$ (7)
@@
$=\prod_{k\in[d]}\prod_{\mathbf{c}\in\langle n\rangle^{k}}w_{k}(\omega\downarrow\mathbf{c})\prod_{k\in[d]}\prod_{\mathbf{c}\in\langle m\rangle^{k}}w_{k}(\omega\downarrow\mathbf{c})\prod_{k\in[d]}\prod_{\mathbf{c}\in\langle n+m\rangle^{k}\setminus\langle n\rangle^{k}\cup\langle\bar{n}\rangle^{k}}\!\!\!\!\!\!\!\!w_{k}(\omega\downarrow\mathbf{c})$
$=w(\omega\downarrow[n])\times w(\omega\downarrow[\bar{n}])\times\prod_{k\in[d]}\prod_{\mathbf{c}\in\langle n+m\rangle^{k}\setminus\langle n\rangle^{k}\cup\langle\bar{n}\rangle^{k}}\!\!\!\!w_{k}(\omega\downarrow\mathbf{c})$
@@
$\prod_{\mathbf{c}\in\langle n+m\rangle^{k}\setminus\langle n\rangle^{k}\cup\langle\bar{n}\rangle^{k}}\!\!\!\!w_{k}(\omega\downarrow\mathbf{c})$ (8)
@@ 6 Domain-Size Generalization
The statement follows from equation (7) in Lemma 2. The upper bound is obtained by replacing the multiplicative weight contribution of each tuple in $\langle n+m\rangle^{k}\setminus\langle n\rangle^{k}\cup\langle\bar{n}\rangle^{k}$ with $w_{k}^{max}$, for all $k\in[d]$. And the lower bound is obtained by replacing the weight contribution of all such tuples with $w_{k}^{min}$. ∎
@@
As the proof of the lower bound follows analogous to the proof of the upper bound, we defer it to the appendix.
@@
Given a Markov Logic Network $\Phi$, then the following inequality holds for all $\omega\in\Omega^{(n)}$:
$\frac{M_{min}}{M_{max}}P_{\Phi}^{(n)}(\omega)\leq P_{\Phi}^{(n+m)}\downarrow[n](\omega)\leq\frac{M_{max}}{M_{min}}P_{\Phi}^{(n)}(\omega)$ (14)
@@
Hence, we have that
$P_{\Phi}^{(n+m)}\downarrow[n](\omega^{\prime})$ $\leq\frac{1}{Z(m)M_{min}C_{n,m}}w(\omega^{\prime})\frac{M_{max}}{Z(n)}\sum_{\omega^{\prime\prime}\in\Omega^{(m)}}\sum_{\begin{subarray}{c}\omega\in\Omega^{(n+m)}\\ \omega\downarrow[n]=\omega^{\prime}\\ \omega\downarrow[\bar{n}]=\omega^{\prime\prime}\end{subarray}}w(\omega^{\prime\prime})$
$=\frac{1}{Z(m)M_{min}C_{n,m}}w(\omega^{\prime})\frac{M_{max}}{Z(n)}\sum_{\omega^{\prime\prime}\in\Omega^{(m)}}C_{n,m}w(\omega^{\prime\prime})$
$=\frac{1}{Z(m)M_{min}C_{n,m}}w(\omega^{\prime})\frac{M_{max}}{Z(n)}C_{n,m}Z(m)$
@@
Note that
$\underset{\omega\in\Omega^{(n)}}{\sum}P^{(n+m)}_{\Phi}\downarrow[n](\omega)\times\log\Delta$
is the expectation value of $\log\Delta$ under the distribution $P^{(n+m)}_{\Phi}\downarrow[n](\omega)$. Since the expectation of a constant is the constant itself, we have that:
$\underset{\omega\in\Omega^{(n)}}{\sum}P^{(n+m)}_{\Phi}\downarrow[n](\omega)\times\log\Delta=\log\Delta$
@@ 7 Experiments
In this section, we evaluate the effect of reducing parameter variance on generalization behavior [^3] Our code is available Online. Proposition 1 bounds $w(\omega)$ w.r.t. the maxima and minima of $w_{k}$. However, for almost all worlds, this bound is loose. This is because, for most worlds, not all $k$-tuples chosen from across the domains will have the extreme weights. Note that in general, our goal is to minimize the impact of the term presented in equation (8). Therefore, it is more effective to reduce the spread between all the weights, rather than merely scaling the upper and the lower bound. Also note that for most MLNs, for some $\omega\in\Omega^{k}$, we will have that $w_{k}(\omega)=1$, i.e., none of the formulas in the MLN will be realized on $\omega$. Thus, in most practical cases, to reduce the spread of the weights $a_{i}$, one should reduce their spread around 00.
Multiple approaches discussed in the literature, directly or indirectly, minimize the parameter variance [8, 17]. We empirically evaluate the effects of three such approaches: L1 regularization, L2 regularization, and Domain-Size Aware Markov Logic Networks (DA-MLNs) [17]. Both L1 and L2 regularization directly work to reduce the spread of the parameters: L1 regularization penalizes the sum of the absolute weight values and L2 regularization penalizes the sum of squared weights. In our setting, we only penalize formulas of arity $>1$, because unary formulas do not affect the connecting term discussed in equation (8).
A DA-MLN is an adaptation of a regular MLN that reduces the variance of the parameters by down-scaling formula weights depending on the domain size of the dataset it should generalize to. In this section, we will call such datasets target sets. A DA-MLN is then given as follows:
@@ 7.1 Datasets
To provide a thorough analysis of the effects of different methods for generalizing across different domain sizes, we use four datasets commonly used in related literature: Friends & Smokers (FS) [25], IMDB [^4] Dataset available on the Alchemy website. [16], WebKB\@footnotemark [16] and Nations\@footnotemark [21].
Friends & Smokers (FS). This synthetic dataset captures information about smoking habits, friendships, and cancer diagnoses of a set of people. The data is created by first randomly selecting 40% of a population to be smokers. Then, 30% of the smokers and 10% of the non-smokers are chosen to suffer from cancer. Lastly, friendships are assigned based on smoking habits, with a $0.8$ probability for friendships between people with the same smoking habit, and a $0.1$ probability of friendships between people with different smoking habits. For our experiments, we generate a target set of size 500.